OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form

OpenStudy (anonymous):

A perpendicular slope, given some slope \[m=\frac{a}{b}\] is:\[m_p=-\frac{b}{a}\]We take the following information and put it into point-slope form:\[(y-y_0)=m(x-x_0)\]\[(y-y_0)=-\frac{1}{4}(x-x_0)\implies (y-2)=-\frac{1}{4}(x-3)\]Solving this in standard form finishes our case.

OpenStudy (anonymous):

Thanks but I still don't understand:/

OpenStudy (anonymous):

Hmm... all right. So, what do we know about the slope of the line we need to draw?

OpenStudy (anonymous):

Well I know I have to use y=mx+ b -3=(-1/4)(2)+b -3=-1/2+b -5/2 giving me... y=-1/4x-5/2 Then Ax+By=C y=-1/4x -5/2 y+1/4x=-5/2 4y+x=-10 Is anything wrong in my work??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Sorry, I was answering another question. But, yes, you are correct, since the slope is -1/4 and (2, -3) is a point as: \[2+4(-3)=-10\]

OpenStudy (anonymous):

No worries! Thank you so very much:D

OpenStudy (anonymous):

Sure thing.